OpenStudy (anonymous):

if I got 15 out of 23 what percentage would that be

OpenStudy (dmezzullo):

like u got 15 outof 23 right?

OpenStudy (anonymous):

yes

OpenStudy (dmezzullo):

ok so what my teacher does is we minus the number wrong from the total number of possible answers

OpenStudy (dmezzullo):

so 23-7 would be 19

OpenStudy (dmezzullo):

then divide 19 into 23

OpenStudy (dmezzullo):

and theres ur percentage

OpenStudy (dmezzullo):

Did da halp any?

OpenStudy (anonymous):

I actually got 8 wrong

OpenStudy (dmezzullo):

ok then 18 divided by 23

OpenStudy (dmezzullo):

Thats right isn't it?

OpenStudy (anonymous):

23-8=15

OpenStudy (dmezzullo):

ok sorry its late my brain aint workin

OpenStudy (dmezzullo):

then divide 15 into 23

OpenStudy (anonymous):

so it would be 65%

OpenStudy (dmezzullo):

yes

geerky42 (geerky42):

\[\text{%} = \left(\dfrac{15}{23}\right) \times100\]

OpenStudy (dmezzullo):

Sorry my brain don't work to well with math at night

OpenStudy (anonymous):

thanks everyone that really helped

geerky42 (geerky42):

Glad we helped.